OpenStudy (anonymous):

find the slope of the tangent to the fnction 4tan^(2)x at the point x=pi/4

OpenStudy (anonymous):

i am getting 16 as the slope am i wrong ??

OpenStudy (anonymous):

i toke the dervative and i got 8tan(pi/4)secx(pi/4)^(2)

OpenStudy (jdoe0001):

looks good, yes

OpenStudy (jdoe0001):

\(\bf sec^2\left(\frac{\pi}{4}\right)\implies sec^2\left(\frac{2}{\sqrt{2}}\right)\implies \cfrac{4}{2}\implies 2\) and the \(\bf tan\left(\frac{\pi}{4}\right)=1\)

OpenStudy (anonymous):

so i am right??

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

i did it so quickly i felt it was wrong

OpenStudy (jdoe0001):

\(\bf sec^2\left({\color{blue}{ \frac{\pi}{4}}}\right)\implies sec^2\left(\frac{2}{\sqrt{2}}\right)\implies \cfrac{4}{2}\implies 2 \\ \quad \\ tan\left({\color{blue}{ \frac{\pi}{4}}}\right)\implies 1 \\ \quad \\ \quad \\ ------------------------\\ \cfrac{c}{dx}[4tan^2(x)]\implies 4\cdot 2tan(x)\cdot sec^2(x) \\ \quad \\ 4\cdot 2tan\left({\color{blue}{ \frac{\pi}{4}}}\right)\cdot sec^2\left({\color{blue}{ \frac{\pi}{4}}}\right)\implies 4\cdot 2\cdot 1\cdot 2\)

OpenStudy (jdoe0001):

hmm ahemm \(\bf sec^2\left({\color{blue}{ \frac{\pi}{4}}}\right)\implies sec^2\left(\frac{2}{\sqrt{2}}\right)\implies \cfrac{4}{2}\implies 2 \\ \quad \\ tan\left({\color{blue}{ \frac{\pi}{4}}}\right)\implies 1 \\ \quad \\ \quad \\ ------------------------\\ \cfrac{d}{dx}[4tan^2(x)]\implies 4\cdot 2tan(x)\cdot sec^2(x) \\ \quad \\ 4\cdot 2tan\left({\color{blue}{ \frac{\pi}{4}}}\right)\cdot sec^2\left({\color{blue}{ \frac{\pi}{4}}}\right)\implies 4\cdot 2\cdot 1\cdot 2\)

OpenStudy (anonymous):

my method is also correct right ? ?

OpenStudy (jdoe0001):

well, I assume you used that =) but it's 16, yes